OpenStudy (anonymous):

Given the function y = 5x – 7, how does the y-value change as x increases by 1? A. The y-value decreases by 7. B. The y-value decreases by 2. C. The y-value increases by 12. D. The y-value increases by 5.

OpenStudy (p0sitr0n):

y'=5x-7 y=5(x+1)-7 y=5x+5-7 y=y'+5 increases by 5, D